# Radiative lifetime of the state in RaF with relevance to laser cooling

### v1

@@ Abstract
The radiative lifetime of the () state in radium monofluoride (RaF) is measured to be 35(1) ns. The lifetime of this state is of relevance to the laser cooling of RaF via the optically closed transition, which is an advantageous aspect of the molecule for its promise as a probe for new physics. The radiative decay rate s is extracted using the lifetime, which determines the natural linewidth of 4.6(3) MHz and the maximum photon scattering rate of 4.1(3)10 s of the laser-cooling transition. RaF is thus found to have a comparable photon-scattering rate with other laser-cooled molecules, while thanks to its highly diagonal Franck-Condon matrix it is expected to scatter an order of magnitude more photons when using 3 cooling lasers before it decays to a dark state. The lifetime measurement in RaF is benchmarked by measuring the lifetime of the state in Fr to be 83(3) ns, in agreement with literature.
@@ I Introduction
The properties of the lowest-lying state in alkaline-earth monofluorides and other molecules of similar electronic structure are of central importance for their suitability for direct laser cooling and trapping. The diagonal Franck-Condon matrix and closed optical cycle between the ground state and the lowest-lying state of these molecules has enabled the direct laser cooling of molecules to mK temperatures or lower, such as CaF [1], SrF [2], YbF [3], YO [4], CaOH [5], SrOH [6], and the symmetric-top CaOCH [7].
The radiative lifetime of the state is key for implementing transversal and longitudinal laser slowing based on the optically closed transition. One-dimensional laser slowing has proven to be very effective in decelerating the molecules to the capture velocity of a magneto-optical trap [8], while simultaneously compressing and collimating the radially expanding molecular beam, leading to a higher density of trapped molecules [3, 4].
As the heaviest alkaline-earth monofluoride, RaF is a radioactive molecule whose longest-lived isotopologue, RaF, has a half-life of 1,600 years due to the decay of the radium nucleus. Thanks to the high atomic number of Ra (), the ground state of RaF has a very strong effective electric field along the internuclear axis at full polarization ( GV/cm) [9], which provides high sensitivity to the electric dipole moment of the electron [10]. Furthermore, the octupole-deformed shape of the Ra nuclei [11] enhances the sensitivity of precision experiments with RaF to a number of symmetry-violating nuclear and hadronic properties [12], such as the nuclear Schiff moment [10, 13]. Moreover, like the other alkaline-earth monofluorides, the electronic structure of RaF is suitable for laser cooling, as originally proposed by ab initio quantum chemistry and supported by recent experiments [14]. Laser cooling has the potential of increasing the interaction time with the molecules by orders of magnitude, thus improving the precision of experimental searches for new physics. Other radioactive molecules, which might also possess a similar laser-coolable electronic structure but whose spectroscopy so far remains challenging and out of reach, are also envisioned for precision tests of fundamental physics [15]. Understanding the structure of RaF, for which experimental study is possible, can also be of general interest to assess the potential of radioactive molecules for fundamental studies.
@@ II Methods
The measurements were obtained via delayed multi-step ionization with the Collinear Resonance Ionization Spectroscopy (CRIS) experiment using beams of RaF and Fr produced at the ISOLDE radioactive ion beam facility at CERN. Details on the production of the radionuclide beams and the CRIS technique can be found in the Supplemental Material and in Refs. [18, 19, 20].
@@
All lasers in this study were pulsed and broadband, with linewidths GHz. The 753-nm step in RaF (Fig. 1) was produced with the fundamental output of a titanium:sapphire (Ti:Sa) laser, while the 423-nm excitation step in Fr was produced by second-harmonic generation of the same Ti:Sa laser. The 646-nm step in RaF was produced using a dye laser pumped by a 532-nm Nd:YAG laser. The 532-nm and 1064-nm steps were produced by the second and fundamental harmonics of a Nd:YAG laser, respectively. The 646-nm and 532-nm steps in RaF and 1064-nm step in Fr are referred to as the ionization steps, while the 753-nm and 423-nm steps are referred to as the excitation steps in the respective schemes in this work.
@@
The pulses of the laser used for the excitation steps had a FWHM of 38(2) ns, while the pulses of the Nd:YAG lasers used for the ionization steps had a FWHM of 15(1) ns. The jitter of the excitation step was measured at 4 ns and that of the ionization steps at 3 ns, for a total relative jitter of 7 ns between excitation and ionization steps in both schemes. As the 646-nm step in the RaF scheme was provided by a pulsed dye laser pumped by an identical Nd:YAG laser as the one used for the 532-nm ionization step, the FWHM and jitter of the second and third steps were measured to be the same, as expected.
@@
To determine the lifetime of the state in RaF and the state in Fr, multiple measurement cycles were performed for various timing arrangements between the ejection of the molecular/atomic beam from the gas-filled linear Paul trap of the ISOLDE facility [21] (bunch temporal spread of 5 s, longitudinal spatial spread of 0.8 m) and the timing of the laser pulses. Firstly, an optimal setting for the Paul trap ejection timing was determined based on the maximum laser-molecule/atom spatial and temporal overlap, and a lifetime measurement was taken for these settings, by moving the timing of the ionization steps forward in time while keeping the timing of the excitation step fixed. One lifetime measurement (denoted as Cycle 1 for both RaF and Fr) was performed for the optimal ejection timing in terms of laser-molecule overlap, and lifetime measurements were also performed for ejection timings 0.5 s earlier and 0.5 s later compared to the optimal settings. One lifetime measurement was also performed at the optimal ejection timing, but instead of moving the ionization steps forward in time, as in the other measurements, the excitation step timing was moved backwards while keeping the timing of the ionization steps fixed. Only for the RaF experiment, an additional measurement (Cycle 4 in Fig. 2c) had been performed in a separate experimental campaign 2 years prior to the rest of the cycles; details about this measurement cycle can be found in the Supplemental Material.
The sum of the pulse FWHM and jitter for the excitation and ionization steps was approximately 60 ns. As a result, for the first few data points in the decay curve shown in Fig. 2b, the excitation and ionization steps were at least partly overlapped in time. As a result, fitting the full curve with ionization delay values down to zero-delay would require modeling the transient population of the state, where stimulated excitation and de-excitation as well as radiative decay are all involved. To extract the radiative lifetime of the state, the analysis becomes significantly simplified if only population loss via radiative decay is considered, as the evolution of the count rate as a function of ionization delay will follow the natural decay law:
where is the population of molecules in the state as a function of time, the initial population in the state at zero delay (after the startpoint, as discussed below), the radiative lifetime of the state, and is a constant representing the ion background on the detector that stems from non-resonant ionization from the ground state due to the high-power lasers. In the case of delayed ionization, can be interpreted as the delay between the excitation step to the state and the subsequent ionization steps.
Eq. 1 allows identifying the subset of data points in Fig. 2b where the excitation and ionization steps are no longer overlapped in time, and thus the count rate as a function of ionization delay can be fully modeled as population loss by the radiative decay of the state. The data subset where only radiative decay takes place was thus identified by fitting gradually smaller subsets (removing the lowest-delay point in each iteration) of the decay curve and tracing the convergence of the reduced of the fit at a minimum value. For the subsets that include data points where the excitation and ionization steps are partly overlapped, Eq. 1 does not suffice to fully describe the dynamics of the system, and therefore the reduced will be greater. Once the subsets no longer include points with partly overlapped laser pulses, the reduced is expected to plateau at a minimum value. This analysis is shown in Fig. 2a. While the plateau is reached at 40 ns, the startpoint for the final analysis is chosen at 60 ns (as illustrated in Fig. 2b), to also consider the sum of the FWHM of the excitation and ionization lasers, as well as their jitter (see Supplemental Material for details). The lifetime extracted from fitting the decay curve with different starting delays for the fit, from 40 to 70 ns, resulted in consistent lifetime values. The same analysis procedure was followed for the Fr measurements.
The final value and its error for the lifetime in both RaF (Fig. 2c) and Fr (Fig. 3) were extracted by considering the lifetime of all the different measurement cycles, thus incorporating measurements covering different experimental conditions.
@@
To test the accuracy of the experimental approach and analysis procedure, the lifetime of the state in Fr was measured and compared with a literature value from Aubin et al. [17]. The weighted mean value of the independent measurements presented in Fig. 3 for Fr is 83(3) ns, consistent with the literature value of 83.5(15) ns [17]. Fitting the values for the different cycles (Fig. 3) with a constant parameter gives a lifetime of 82(4) ns that is also consistent within 1. Following the same approach, the weighted mean result for the lifetime of the state in RaF is 35(1) ns.
@@ III Results and Discussion
The lifetime of the state in RaF was measured to be ns. The radiative lifetime and excitation energy of the () states across alkaline-earth monofluorides (and the homoelectronic YbF) are shown in Fig. 4. A trend of gradually increasing lifetime for increasing molecular mass is evident, which appears correlated with a decrease in the excitation energy of the states. While the lifetime in RaF matches the observed trend well, BaF appears to be an outlier with a lifetime more than twice as large as expected by the observed trend. This is accompanied by a local minimum in the excitation energy. The distinct properties of this state in BaF are a result of perturbations from the neighboring states, which lie energetically lower than the state in BaF [22], contrary to the other alkaline-earth monofluorides.
The excitation energy with respect to the ground state is of importance for the laser-cooling scheme, as the wavelength of the cooling transition determines the single-photon recoil velocity, , where is the mass of the molecule, , and is the transition wavelength. The trends in Fig. 4 imply that laser-cooling efficiency decreases for heavier alkaline-earth monofluorides, as the molecular mass and transition wavelength increase and thus the single-photon recoil velocity decreases, while the increasing radiative lifetime of the upper state leads to a reduced photon scattering rate. For the laser-cooling transition in RaF, the recoil velocity is equal to 0.22 cm s.
@@
| Radiative decay rate (s) | 2.9(2)10 |
| Maximum scattering rate (s) | 4.1(3)10 |
@@
| Recoil velocity (cm s) | 0.22 |
| Doppler limit (K) | 111(10) |
| Scattering acceleration (km s) | 9.0(8) |
@@
Based on the lifetime ns, the radiative decay rate s can be extracted. is a key quantity for a number of spectroscopic properties, such as the natural linewidth () of the transition, which is found to be equal to 4.6(3) MHz, and the maximum photon scattering rate of the transition. For the laser-cooling scheme proposed in Ref. [14], can be approximated for realistic conditions of a magneto-optical trap as , as summarized in Ref. [31], which in this case is equal to s. With this scattering rate, the optically closed transition in RaF could scatter 100,000 photons in 24.4(2) ms, assuming identical for the states of , which are involved in the laser-cooling scheme [14]. Considering the photon recoil velocity of 0.22 cm s, the maximum scattering rate implies a photon scattering acceleration of km s. These results are summarized in Table 1.
@@
Fig. 5 compares the expected performance of laser cooling for RaF and stable diatomic molecules that have been successfully cooled, all featuring a laser-cooling cycle based on optically closed - transitions. Due to its highly diagonal Franck-Condon matrix for the - transition [14], which far exceeds the diagonality of the transition in the other molecules, the - transition in RaF can scatter an order of magnitude more photons when using 3 cooling lasers before a significant fraction of the molecules is lost to higher vibrational states, while potentially scattering up to 10 photons using only 2 lasers that address the vibrational states, as shown in Fig. 5a.
@@
Although the radiative decay rate is the smallest for RaF, the time required to scatter 10 photons, which can also serve as a measure of the time required to slow a molecular beam down to the capture velocity of a magneto-optical trap, is comparable to that for other molecules (Fig. 5b). Molecular beams produced by cryogenic buffer-gas cells, which most often serve as the source in molecular trapping setups, are emitted with forward velocities that appear to be weakly dependent on the molecular mass of the beam [34, 35]. Fig. 5c compares the distance required to slow a molecular beam down to the Doppler limit using counter-propagating laser cooling. For a forward velocity of 50 m/s, which is achievable with a two-stage buffer-gas cell at 1 K, the difference in slowing RaF compared to the lighter molecules becomes insignificant for a beamline of realistic dimensions.
@@
As a result, the diagonality of the transition in RaF, which could allow effective laser cooling with fewer lasers, is accompanied by a scattering rate that is comparable to that of other laser-cooled molecules. It should be noted that a more efficient cooling scheme exists for CaF that utilizes the exceptionally diagonal - transition, but this approach is not applicable to RaF.
@@
Beyond RaF, other radium-containing molecules have also been proposed as promising candidates for searches of new physics based on calculations of a laser-coolable electronic structure. Especially appealing candidates are polyatomic molecular ions, as they combine the long coherence time achievable in ion traps, which is crucial for the achievable precision of an experimental search for new physics [36], with the enhanced ability to fully polarize polyatomic molecules [37]. Several radium-containing polyatomic cations have been theoretically investigated for their suitability for direct laser cooling [38, 39], and while they are expected to possess a highly diagonal Franck-Condon matrix for the transition of relevance, the calculated excited-state lifetime of the investigated species is in the range of a few s [39]. On the contrary, radium-containing molecules that are expected to be laser-coolable as neutral species appear to have lifetimes in the ns range, such as ns for RaOH [40], making them similarly promising candidates as RaF for laser cooling and trapping. Therefore, experimental methods for the measurement of symmetry-violating moments based on laser-slowed and -cooled neutral molecules [41] are expected to be critical for progress in designing sensitive searches for new physics with radium-containing molecules.

### v2

@@ Abstract
The radiative lifetime of the () state in radium monofluoride (RaF) is measured to be 35(1) ns. The lifetime of this state and the related decay rate s-1 are of relevance to the laser cooling of RaF via the optically closed transition, which makes the molecule a promising probe to search for new physics. RaF is found to have a comparable photon-scattering rate to homoelectronic laser-coolable molecules. Thanks to its highly diagonal Franck-Condon matrix, it is expected to scatter an order of magnitude more photons than other molecules when using just 3 cooling lasers, before it decays to a dark state. The lifetime measurement in RaF is benchmarked by measuring the lifetime of the state in Fr to be 83(3) ns, in agreement with literature.
@@ I Introduction
The properties of the lowest-lying state in alkaline-earth monofluorides and other homoelectronic molecules are of central importance for their suitability for direct laser cooling and trapping. The diagonal Franck-Condon matrix and closed optical cycle between the ground state and the lowest-lying state of these molecules has enabled the direct laser cooling of molecules to mK temperatures or lower, such as CaF [1], SrF [2], YbF [3], YO [4], CaOH [5], SrOH [6], and the symmetric-top CaOCH3 [7]. The short radiative lifetime of the state is key for efficient laser cooling based on the optically closed transition, and decelerating the molecules to the capture velocity of a magneto-optical trap [8].
The homoelectronic and radioactive RaF has been proposed as a sensitive probe for searches of parity or time-reversal () violating properties, such as the electric dipole moment of the electron (eEDM). Extending laser cooling and trapping to RaF would enable highly sensitive searches for new physics by significantly increasing the interrogation time of the RaF molecules.
Thanks to the high atomic number of Ra (), the ground state of RaF, whose longest-lived isotopologue 226Ra19F has a half-life of 1,600 years, is highly sensitive to the eEDM [9, 10]. Furthermore, the octupole-deformed shape of the 222-226Ra nuclei [11] also enhances the sensitivity of RaF to a number of -violating nuclear properties [12], such as the Schiff moment [9, 13]. Combined with the potential to achieve exceptionally long coherence times thanks to laser cooling, RaF is a promising system to search for new physics with potentially unprecedented sensitivity.
@@ II Methods
The measurements were obtained via delayed multi-step ionization with the Collinear Resonance Ionization Spectroscopy (CRIS) experiment using beams of 226Ra19F+ and 221Fr+ produced at the ISOLDE radioactive ion beam facility at CERN. The technique and the employed laser schemes are shown schematically in Fig. 1, and further experimental details can be found in the Supplemental Material and in Refs. [16, 17, 18].
Fast ions (40 keV kinetic energy) of 226Ra19F+ and 221Fr+, released as a pulsed beam from a linear Paul trap (100-Hz repetition, 5-s temporal spread, 0.8-m longitudinal spatial spread), were delivered to the CRIS experiment and were neutralized via collisions with a vapor of sodium atoms in a charge-exchange cell, while the remaining ions were deflected onto a beam dump. The neutralized bunches entered the interaction region, where they were temporally and spatially overlapped in a collinear geometry with pulsed laser beams (see Fig. 1a) that step-wise excited the molecules/atoms from the ground state to above the ionization potential, as per the laser schemes shown in Fig. 1b,c. The resonantly re-ionized molecules were deflected onto a MagneToF single-ion detector, while the residual neutral beam was discarded onto a beam dump.
@@
All lasers in this study were pulsed and broadband, with linewidths GHz. The 753-nm step in RaF was produced with the fundamental output of a titanium:sapphire (Ti:Sa) laser, while the 423-nm excitation step in Fr was produced by second-harmonic generation of the same Ti:Sa laser. The 646-nm step in RaF was produced using a dye laser pumped by a 532-nm Nd:YAG laser. The 532-nm and 1064-nm steps were produced by the second and fundamental harmonics of a Nd:YAG laser, respectively. The 646-nm and 532-nm steps in RaF and 1064-nm step in Fr are referred to as the ionization steps, while the 753-nm and 423-nm steps are referred to as the excitation steps in the respective schemes in this work.
@@
The pulses of the laser used for the excitation steps had a full width at half-maximum (FWHM) of 38(2) ns, while the pulses of the Nd:YAG lasers used for the ionization steps had a FWHM of 15(1) ns. The shot-to-shot jitter of the excitation step was measured to be 4 ns and that of the ionization steps to be 3 ns, for a total relative shot-to-shot jitter of 5 ns (added in quadrature) between excitation and ionization steps in both schemes (details in the Supplemental Material). As the 646-nm step in the RaF scheme was provided by a pulsed dye laser pumped by an identical Nd:YAG laser as the one used for the 532-nm ionization step, the FWHM and jitter of the second and third steps were measured to be the same, as expected.
@@
To determine the lifetime of the state in RaF and the state in Fr, multiple measurement cycles were performed for various timing arrangements between the ejection of the molecular/atomic beam from the gas-filled linear Paul trap of the ISOLDE facility [19] and the timing of the laser pulses. Firstly, an optimal setting for the Paul trap ejection timing was determined based on the maximum laser-molecule/atom spatial and temporal overlap, and a lifetime measurement was taken for these settings, by moving the timing of the ionization steps forward in time while keeping the timing of the excitation step fixed. Lifetime measurements (denoted as Cycle 1 for both RaF and Fr) were performed for this optimal ejection timing, as well as for ejection timings 0.5 s earlier and 0.5 s later compared to the optimal for Fr (1-s steps for RaF). One lifetime measurement was also performed at the optimal ejection timing, but moving the excitation step timing backwards while keeping the timing of the ionization steps fixed. Only for the RaF experiment, an additional measurement (Cycle 4 in Fig. 2c) had been performed in a separate proof-of-principle experimental campaign 2 years prior to the rest of the cycles; details about this measurement cycle can be found in the Supplemental Material.
Fig. 2a shows an example of the measured ion count rate as a function of the delay time between the excitation step and the subsequent ionization steps (all decay curves are shown in the Supplemental Material). For delay times larger than the sum of the pulse widths (FWHM) and the relative jitter, the data represents the radiative decay from the excited state, and the curve can then be fitted with the natural decay law:
In Eq. 1, is the population of molecules in the state as a function of the delay, is the population at delay time , is the radiative lifetime of the state, and is a constant representing the ion background on the detector that stems from non-resonant ionization. In order to exclude data points where the radiative decay law is not valid (at times when there is an overlap of the lasers), the first 60 ns were omitted when fitting the decay curve, shown as the ”laser overlap regime” in Fig. 2a. This span of 60 ns is approximately equal to the sum of the FWHM and relative jitter of the lasers (see Supplemental Material for details).
The laser overlap regime was confirmed as shown in Fig. 2b, where the delay time at which the reduced saturates is verified, which marks the end of the laser overlap regime. While saturation already start from 40 ns, all decay curves were analyzed with a 60-ns start point to exclude data with partial laser overlap. Fitting the decay curves with start points between 40 to 70 ns resulted in consistent lifetime values. The same procedure was followed for Fr.
A 5-ns error was adopted for the laser delay time, which corresponds to the combined shot-to-shot laser jitter of the excitation and ionization lasers (see Supplemental Material). Repeating the fit procedure with a smaller x-error yielded a reduced progressively closer to 1, while the lifetime result remained unchanged. The uncertainties on the final lifetime values were not scaled with the respective reduced of each fit, as this would reduce the uncertainty, and it is preferred to remain conservative in the result.
The fit residuals for the decay curve are shown in Fig. 2a. The fraction of data points whose deviation from the fit exceeds the -error is consistent with the 1 confidence interval. A potential oscillatory trend in the residuals, not seen in all data sets (see Fig. 4 in Supplementary Material), falls within the confidence interval.
In Fig. 2c, the final results from the lifetime analysis for RaF are presented, extracted from the different measurement cycles that cover different experimental conditions. The weighted mean value of the independent measurements is found to be 35(1) ns.
@@
To test the accuracy of the experimental approach and analysis procedure, the lifetime of the state in Fr was measured using the same approach, and compared with a literature value from Aubin et al. [15]. The results are presented in Fig. 3, leading to a weighted mean value of 83(3) ns, consistent with the literature value of 83.5(15) ns [15].
@@ III Results and Discussion
The lifetime of the state in RaF was measured to be ns. The radiative lifetime and excitation energy of the () states across alkaline-earth monofluorides (and the homoelectronic YbF) are shown in Fig. 4. A trend of gradually increasing lifetime for increasing molecular mass is evident, which appears correlated with a decrease in the excitation energy of the states.
The excitation energy is of importance for the laser-cooling scheme, as it determines the single-photon recoil velocity, , where is the mass of the molecule, , and is the transition wavelength. The trends in Fig. 4 imply that laser-cooling efficiency decreases for heavier alkaline-earth monofluorides, as the molecular mass and transition wavelength increase, while the increasing lifetime of the upper state leads to a reduced photon scattering rate. For the laser-cooling transition in RaF, the recoil velocity is equal to 0.22 cm s-1.
@@
Radiative decay rate (s-1) | 2.86(8)107 |
Maximum scattering rate (s-1) | 4.1(1)106 |
@@
Recoil velocity (cm s-1) | 0.22 |
Doppler limit (K) | 109(3) |
Scattering acceleration (km s-2) | 9.0(3) |
Based on the lifetime ns, the radiative decay rate s-1 can be extracted. is a key quantity for a number of spectroscopic properties, such as the natural linewidth () of the transition, which is found to be equal to 4.6(1) MHz, and the maximum photon scattering rate of the transition.
@@
For the laser-cooling scheme proposed in Ref. [29], can be approximated for optimistic conditions of a magneto-optical trap as , based on a set of rate equations as summarized in Ref. [30], which in this case is equal to s-1. With this scattering rate, the optically closed transition in RaF could scatter 100,000 photons in 24.4(7) ms, assuming identical for the states of [29]. It is noted that the accuracy of this expression for is limited by the validity of the rate model in a practical setting [30], and these results thus represent optimistic values. It is possible that experimental measurements of the scattering rate could be lower than this value.
Lastly, considering the photon recoil velocity of 0.22 cm s-1, the maximum scattering rate implies a photon scattering acceleration of km s-2. These results are summarized in Table 1.
@@
Fig. 5 compares, based on the lifetime of the state and the Franck-Condon factors, the expected performance of laser cooling for RaF and stable diatomic molecules that have been successfully cooled, all featuring a laser-cooling cycle based on optically closed - transitions. Due to its highly diagonal Franck-Condon matrix for the - transition [29], which far exceeds the diagonality of the transition in the other molecules, the - transition in RaF can scatter an order of magnitude more photons when using 3 cooling lasers before a significant fraction of the molecules is lost to higher vibrational states, while potentially scattering up to 105 photons using only 2 lasers that address the vibrational states, as shown in Fig. 5a.
@@
Although the radiative decay rate is the smallest for RaF, the time required to scatter 105 photons, which can also serve as a measure of the time required to slow a molecular beam down to the capture velocity of a magneto-optical trap, is comparable to that for other molecules (Fig. 5b). Molecular beams produced by cryogenic buffer-gas cells, which most often serve as the source in molecular trapping setups, are emitted with forward velocities that appear to be weakly dependent on the molecular mass of the beam [33, 34]. Fig. 5c compares the distance required to slow a molecular beam down to the Doppler limit using counter-propagating laser cooling. For a forward velocity of 50 m/s, which is achievable with a two-stage buffer-gas cell at 1 K, the difference in slowing RaF compared to the lighter molecules becomes insignificant for a beamline of realistic dimensions.
@@
As a result, the diagonality of the transition in RaF, which could allow effective laser cooling with fewer lasers, is accompanied by a scattering rate that is comparable to that of other laser-cooled molecules. A more conclusive discussion on the number of photons scattered by RaF per cooling laser would require direct measurements of vibrational branching ratios in the future.
@@
Beyond RaF, other radium-containing molecules have also been proposed for searches of new physics based on calculations of a laser-coolable electronic structure. Neutral polyatomic radium-containing molecules appear to have lifetimes in the ns range, such as ns for RaOH [35], making them similarly promising candidates as RaF for laser cooling and trapping. Therefore, experimental methods for the measurement of symmetry-violating moments based on laser-slowed and -cooled neutral molecules [36] are expected to be critical for progress in designing sensitive searches for new physics with radium-containing molecules.